OpenStudy (anonymous):

Convert the polar equation to rectangular coordinates. r = 6 cos θ

OpenStudy (anonymous):

I got is this correct? x+y=6

OpenStudy (experimentx):

put \[ r = \sqrt{(x^2+y^2)}\] and \[ \cos\theta = \frac{x}{\sqrt{(x^2+y^2)}}\] rest of the work is to simplify

OpenStudy (anonymous):

Im not seeing it

OpenStudy (anonymous):

did the problem over now i get (x-3)^2 + y^2 = 6

OpenStudy (experimentx):

looks like you got a circle ... though i am not quite sure myself on polar coordinates.

OpenStudy (experimentx):

looks like thats right http://www.wolframalpha.com/input/?i=r+%3D+6+cos+%CE%B8

Directrix (directrix):

>>I got is this correct? x+y=6 ---- No. r = 6 cos (theta) is a circle. x + y = 6 is a line. r = 6 cos (theta) r = 6 (x/r) Note: x/r is the ratio of the adjacent leg to the hypotenuse (radius) r/1 = (6x)/r r^2 = 6x x^2 + y^2 = 6x y^2 = 6x - x^2

OpenStudy (anonymous):

i got x^2=(y-3)^2=1

OpenStudy (anonymous):

@experimentX